OpenStudy (anonymous):

Determine if the vectors are orthogonal, parallel or neither v = <3i - 5j>, w=< -9i +10j

OpenStudy (anonymous):

What's the dot product of the two vectors?

OpenStudy (anonymous):

That's all I have

OpenStudy (jdoe0001):

well.. do you know how to find the dot product of two vectors? I assume this has already been covered depending on the dot product of the two vectors, you can see if they're such if 0 is the dot product, then they are orthogonal and they're parallel if their coordinates are multiplies of each other

OpenStudy (anonymous):

when I did the problem and the answer I got wasn't zero nor was it a multiple of each other. so I assumed it was neither

OpenStudy (jdoe0001):

yeap, then is neither

OpenStudy (anonymous):

is that the same answer you got?

OpenStudy (anonymous):

A non-zero dot product only guarantees that the two vectors aren't orthogonal. To check if they're parallel or neither, you have to use a handy formula: \[a\cdot b=|a||b|\cos\theta\] where \(a\) and \(b\) are the given vectors, \(|a|\) and \(|b|\) are their respective magnitudes, and \(\theta\) is the angle between them.

OpenStudy (anonymous):

If the angle is 0 or 180 degrees, they're parallel. Otherwise, neither.

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (jdoe0001):

on the other hand, they're not multiple of each other =)

OpenStudy (anonymous):

thanks Guys

OpenStudy (anonymous):

yw